OpenStudy (anonymous):

Can someone help this? The derivative of y = (-x)/(1-x^2)^0.5 is ?

OpenStudy (anonymous):

do you have the selected answer?

OpenStudy (anonymous):

love derivatives. so you are using a couple different things here (also if you want a very nice step by step try wolframalpha, i get sloppy sometimes). first you need to use the quotient rule which is: (f(x)/g(x))'= [f'(x)g(x)-g'(x)f(x)]/g(x)^2 so knowing that we are eventually going to be putting it in that form, lets first find the derivative of the top. this one isn't too bad, since the derivative of x is 1. so factoring in the - sign we get: -x'=-1 now lets look at the bottom part. since this is a little more complicated, we are going to use the chain rule. so we need to take the derivative of the outside times the derivative of the inside: (1-x^2)^(1/2)'=1/2{[1-x^2]^(-1/2)}(-2x) reducing that we get: -x/(1-x^2)^(1/2) now lets put it all together: -(1-x^2)^(1/2)-(-x)(-x/(1-x^2)^(1/2)) / 1-x^2

OpenStudy (anonymous):

i feel like i should go to a book signing now.

OpenStudy (anonymous):

hmm my answer is [-1+(x^2)(1-x^2)^-(3/2) ]

OpenStudy (anonymous):

true story, forgot the exponent on the bottom was negative. uber fail

OpenStudy (anonymous):

actually also wrong, i just wolframed that pellet. it gave me: -1/(1-x^2)^(3/2)

OpenStudy (anonymous):

try to cancel the denominator by factoring out the 1-x^2

OpenStudy (anonymous):

ah ok good

OpenStudy (anonymous):

i got (-1-x^2)/(1-x^2)^(3/2)

OpenStudy (anonymous):

you are very close try again..continue solving til you get y=-1/(1-x^2)^(3/2)

OpenStudy (anonymous):

can you show me how to get it?i still cant get it.

OpenStudy (anonymous):

ryt, did you use the derivative of the quotient formula? use that

OpenStudy (anonymous):

d(u/v)=(vdu-udv)/v^2

OpenStudy (anonymous):

yea i used it but cant get that answer.

OpenStudy (anonymous):

why is it 1 -2x? i thought it will be -2x ?

OpenStudy (anonymous):

dy/dx = d[(-x)/(1-x^2)^0.5]/dx =((1-x^2)^1/2)(-1)-(-x)(1/2)(1-x^2)^-1/2(-2x)]/(1-x^2)

OpenStudy (anonymous):

yes -2x, sorry wrong typing lol...

OpenStudy (anonymous):

oh okay. so the answer is (-1-x^2)/(1-x^2) ?

OpenStudy (anonymous):

=(-(1-x^2)^1/2)-(x^2)((1-x^2)^-1/2]/(1-x^2)

OpenStudy (anonymous):

no the answer is dy/dx =-1/(1-x^2)^(3/2)

OpenStudy (anonymous):

how you get that from the previous working?

OpenStudy (anonymous):

=(-(1-x^2)^1/2)-(x^2)((1-x^2)^-1/2]/(1-x^2) , factoring out (1-x^2) =(1-x^2)[-(1-x^2)^1/2 -x^2 (1-x^2)^-3/2]/(1-x^2) =[-(1-x^2)^1/2 -x^2 (1-x^2)^-3/2] =-1/(1-x^2) -x^2 /(1-x^2)^3/2 =-(1-x^2)-x^2 /((1-x^2)^3/2 thats it.. =-1/(1-x^2)^3/2 ans

OpenStudy (anonymous):

did you get it ryt?

OpenStudy (anonymous):

how to do you get the answer from the pervious one?(5th and 6th line of working)

OpenStudy (anonymous):

hi ryt sorry my pc hanged up, i didnt get your question on time

OpenStudy (anonymous):

oh its okay:) i get it :D thanks!!